« Status: 21FEB07 (Wednesday - morning) | Main | Status: 21FEB07 (Wednesday - sunset) »

Status: 21FEB07 (Wednesday - midday)

IOP-6 is in progress, but the storm is weaker and considerably slower than predicted yesterday. The onset of precip in the ARB is now not expected before about 4 pm today. Nevertheless, today's models are still predicting 1.5 to 2.5 inches of precipitation (liq equiv) from this system by Friday afternoon. Snow level should drop from 5 kft today to 1.5 kft on Friday. Sloughhouse and Reno soundings have been going up at 4-hr intervals since 12 UTC this morning, and will continue. SMART-R is set to go, but will hold off on data collection until the precip approaches the foothills. No rainwater chemistry sampling is being conducted in this storm. IOP-6 operations are expected to continue into Friday.

Problems (besides a sluggish storm):
There is a computer problem for the ESRL site at Cazadero in the Coast Range; manual intervention is being considered to solve that. The ESRL hi-rez model package suffered a major outage when a new compiler crashed on Monday and shut the system off since then. Availability of SHS soundings on the ALPS workstations in Monterey and CNRFC is still an issue. GSD is working to solve these problems.

The outlook for the next storms includes a weak (QPF ~ 0.6 inches) weekend event, followed by a more promising (QPF ~ 2.6 inches) storm on Tues. ECMWF has the second storm looking even stronger.

- Martner


TrackBack URL for this entry: